# Detecting Invariant Manifolds in ReLU-Based RNNs

### v1

@@ H.2 Additional algorithms
In the main text, the algorithm was formulated for the PLRNN. The same procedure as described in Algo. 2 can analogously be applied to a 1-hidden layer PLRNN, called the shallow PLRNN (shPLRNN) [Hess et al., 2023]:
@@
where ${\bm{A}}\in\mathbb{R}^{M\times M}$, ${\bm{W}}_{1}\in\mathbb{R}^{M\times H}$, ${\bm{W}}_{2}\in\mathbb{R}^{H\times M}$, ${\bm{h}}_{2}\in\mathbb{R}^{H}$, ${\bm{h}}_{1}\in\mathbb{R}^{M}$ and $H>M$. The inversion of this map yields
@@
Fallback algorithm
For systems with complex dynamics, such as in Fig. 3B where trajectories jump between disjoint subregions, the primary algorithm may struggle to converge. In such cases, we use a more robust fallback method (Alg. 3) that perturbs seed points along the analytically defined local manifold and iterates $F_{\mathbf{\theta}}$ to generate dense support vectors. As manifolds can re-enter the same subregion in multiple folds, we apply HDBSCAN [McInnes et al., 2017, Ester et al., 1996] to cluster support vectors into distinct segments. Although computationally more demanding, this fallback reliably captures manifolds with discontinuous or folding structures when sequential tracing fails.
Algorithm 3 Finding stable/unstable manifolds: fallback algorithm
[…]
6:  for $v^{u}\in E^{u}$ do
7:   $z_{0}\mathrel{+}=v^{u}\cdot rand()\trianglerightPerturbe$ into subspace
8:  end for
9:  $T^{u}\leftarrow\textit{GetForwardTS}(z_{0})$
10:  $S^{u}\leftarrow S^{u}\cup\{T^{u}\}$
11:end for
[…]
13:  $z_{0}=P$ $\triangleright$ For $N_{2}$ different initialisations
14:  for $v^{s}\in E^{s}$ do
15:   $z_{0}\mathrel{+}=v^{s}\cdot rand()\triangleright$ Perturb into subspace
16:  end for
17:  $T^{s}\leftarrow\textit{}{GetBackwardTS}(z_{0})$
18:  $S^{s}\leftarrow S^{s}\cup\{T^{s}\}$
19:end for
[…]
22:for each $D\in D_{\Omega}$ do
23:  $S^{s}_{\Omega}\leftarrow S^{s}\cap D_{\Omega}$ $\triangleright$ Go through all subregions
24:  $S^{u}_{\Omega}\leftarrow S^{u}\cap D_{\Omega}$
25:  $(C^{s}_{\Omega},C^{u}_{\Omega})\leftarrow\textsc{Fit}((S^{s}_{\Omega},\,S^{u}_{\Omega}))$ $\triangleright$ Cluster points and fit
26:  $\widetilde{S}^{s}\leftarrow\widetilde{S}^{s}\cup\{C^{s}_{\Omega}\}$
27:  $\widetilde{S}^{u}\leftarrow\widetilde{S}^{u}\cup\{C^{u}_{\Omega}\}$
28:end for
29:return $(\widetilde{S}^{s},\,\widetilde{S}^{u})$ $\triangleright$ Piecewise linear manifolds
Algorithm 4 Additional functions

### v3

@@ Appendix C Additional Details on Manifold Construction
Backtracking algorithm and other PLRNN variants
Step 2 in sect. 3.3 outlines the steps for inverting the RNN map $F_{\mathbf{\theta}}$ for PLRNNs of the form Eq. 1, spelled out in detail in Algorithm 2. This can easily be amended for other ReLU-based RNNs. For instance, for a 1-hidden layer PLRNN, called the shallow PLRNN (shPLRNN) [hess], given by
@@
where ${\bm{A}}\in\mathbb{R}^{M\times M}$, ${\bm{W}}_{1}\in\mathbb{R}^{M\times H}$, ${\bm{W}}_{2}\in\mathbb{R}^{H\times M}$, ${\bm{h}}_{2}\in\mathbb{R}^{H}$, ${\bm{h}}_{1}\in\mathbb{R}^{M}$, the inversion of this map yields
@@
Manifold modeling
To generate initial conditions on the local manifold in Algorithm 1, we sample perturbations in the stable eigenspace around a base point $\bm{x}_{b}$ (equal to $\bm{p}$ in the initial linear subregion). To account for anisotropy in the vector field and potential undersampling of slow eigendirections, we first construct an orthonormal basis $E_{\text{ortho}}^{\sigma}$ via the QR decomposition ($E^{\sigma}=E^{\sigma}_{\text{ortho}}\,R$) which preserves the span of the stable eigenspace while avoiding numerical degeneracy. Transforming the linear operator into this basis yields the effective eigenvalues $\tilde{\lambda}_{i}=\left(R\,\mathrm{diag}(\lambda_{1},\ldots,\lambda_{d_{s}})R^{-1}\right)_{ii},$ which govern contraction rates along the orthonormalized directions. Because systems with disparate timescales typically exhibit large variability along fast eigendirections but very limited variability along slow ones, naive sampling leads to poorly conditioned point clouds. We therefore rescale perturbations inversely to the magnitude of the effective eigenvalues via $s_{i}=\frac{1}{|\tilde{\lambda}_{i}|^{\,c}}$, where the hyperparameter $c$ controls how strongly slow directions are emphasized. We then draw $N_{s}$ perturbation coefficients from a Gaussian mixture model (GMM), $\bm{\alpha}\sim\frac{1}{3}\sum_{k=1}^{3}\mathcal{N}\!\left(0,\,\sigma_{k}^{2}I\right),\sigma_{k}\in\{0.01,\,0.1,\,0.5\},$ where the three scales $\sigma_{k}$ ensure multi-scale coverage of the local manifold. Eigenvalue-based rescaling is applied multiplicatively ($\tilde{\alpha}_{i}=s_{i}\,\alpha_{i}$). Finally, the sampled points on the manifold are generated by $\bm{x}=\bm{x}_{b}+E^{\sigma}_{\text{ortho}}\bm{\tilde{\alpha}}$, and points violating ReLU-constraints (i.e., are inconsistent with the active set at $\bm{x}_{b}$) are rejected. This procedure yields a numerically well-conditioned set of initial points distributed across all stable eigendirections and across multiple sampling scales.
In the case of curved manifolds, we used kernel-PCA [scholkopf1998nonlinear] for approximation in Algorithm 1. Specifically, we used an RBF kernel, the squared exponential kernel $K({\bm{x}},{\bm{y}})=\text{exp}({-\frac{1}{2}\norm{{\bm{x}}-{\bm{y}}}^{2}})$, for its generally good approximation properties for smooth manifolds with curvature as described by 4 [hofmann2008kernel]. Since $K({\bm{x}},{\bm{y}})$ decays with Euclidean distance, kernel-PCA with an RBF kernel emphasizes local geometry, which is suitable for manifolds with curvature. Moreover, the RBF kernel is universal and capable of representing smooth curvature. While more specialized kernels (e.g., spectral–periodic or polynomial–exponential) may sound like a theoretically better match, they tend to require careful tuning of system-specific frequencies or decay rates. The RBF kernel, by contrast, offers a robust and broadly applicable alternative that performs reliably without region-dependent parameter adjustments.